Title: BL Lac
Session: Oct 1996 (session 3)
Frequency: 6 cm
Obs time (UT): 296 EU 16:30(22/10)-16:30(23/10)
Obs time (UT): 296 US 16:30(22/10)-16:30(23/10)
EU Mk3 Tapes/St.: 3
To give feedback on this experiment, click here
Westerbork (Tony Foley at Fri Oct 25 03:12:33 EDT 1996) Minor failures
The procedure file used the old head command instead of the
current pass command so that passes 1-8 overwrote each other
on the first tape. Passes 9 and above are OK.
Effelsberg (D.Graham at Tue Oct 29 02:09:12 EST 1996) Minor failures
Snap generated locally using drudg program. For EB, gg28.drg
listed 14 head positions, but only gave positions for 12 .
Passes 25-28 were therefore written at an unknown position on tape,
and probably overwrote good data
MPIfR (Scott Aaron at Tue Sep 23 14:21:53 METDST 1997) Success
MPIfR (Walter Alef at Tue Sep 23 17:43:26 METDST 1997) Minor failures
Due to PI scheduling error lost 4 passes for Eb.